A sound wave of frequency 300 Hz travels with speed 340ms-1 along




the negative x-direction. Each point of the medium moves up and down through a




total distance of 5.0 mm. Determine the necessary parameters to represent the




wave mathematically.

1
Expert's answer
2022-01-19T12:40:13-0500

To represent the wave mathematically, we need the angular frequency and amplitude:


"\omega=2\pi f=600\pi\text{ rad/s},\\
A=0.005\text{ m},\\
x(t)=-A\sin(\omega t)=-0.005\sin(600\pi t)\text{ m}."